OpenStudy (anonymous):

simplify 3[(2+4*2^2)-6(3-1)]

OpenStudy (anonymous):

18

OpenStudy (anonymous):

= 3 [( 2 + 16) - 6*2] = 3(18-12) = 18

OpenStudy (anonymous):

thank you (: